# takeawaybox backgroundcolor=gray!30, shadow=false, innerleftmargin=5pt, innerrightmargin=5pt, roundcorner=10pt, skipabove=12pt

### v3

@@ 7.2.2 What Works for Representation Operationalization?
Are matrices or vectors better concept operators? - Matrix more effective but more costly. Multiple works that propose using matrices instead of vectors as concept operators find that they improve steering effectiveness (Postmus & Abreu, 2024; Rajendran et al., 2024; Pham & Nguyen, 2024). Results from Zou et al. (2023a) are more ambivalent, finding that a matrix slightly outperforms a vector-based concept operator, but is clearly outperformed by a vector that is derived from additional forward passes for every new input. While matrices can be more effective, they can also come at an increased cost (Postmus & Abreu, 2024) or reduced output quality (Rajendran et al., 2024).
Are linear or non-linear concept operators better? - Non-linear. While most methods perform linear operations on the representations, the direct comparison between linear and non-linear methods favors the non-linear ones. Non-linear Spectral Editing of Activations (SEA) outperforms linear SEA (Qiu et al., 2024), and non-linear ITI is more effective than ITI, although this comes at a higher cost in capabilities (Hoscilowicz et al., 2024).
@@ 7.2.3 What works for Representation Control?
Is it more effective to modify weights or activations? - Unclear. This question is still open. Wang et al. (2024a) find that their method for controlling model weights is more effective than intervening on activations while leading to a larger reduction in output quality. However, a broader evaluation of that question is necessary.
@@
Which steering function should be used for modifying activations? - Unclear. Zou et al. (2023a) find that a piece-wise operation is more effective than linear addition. Chu et al. (2024) find that projection outperforms a product, which again is more effective than linear addition. Luo et al. (2024) find that linear addition is more effective than orthogonal projections at a higher cost in capabilities. Linear addition remains the most popular activation steering function despite some other steering functions outperforming it. This warrants further investigation comparing different operations. Krasheninnikov & Krueger (2024) study different activation steering functions in a toy classification setup. They find that more expressive steering functions have a higher performance ceiling and can control more complex tasks, but simpler steering functions are effective with lower amounts of data.
@@
At which token positions can activations be steered most effectively? - Apply steering at every inference step. Overall, it is more effective to continuously apply RepE at every inference step than to only apply it at the first token (Subramani et al., 2022). This is because the steering effect diminishes throughout the generation (Scalena et al., 2024). However, only intervening on the first token can still be very effective (Subramani et al., 2022) and has a lower impact on output quality (Scalena et al., 2024).
@@
Which component of the transformer should be steered? - Unclear. RepE often modifies activations in the residual stream, the attention or the MLP block, but it is unclear which component is most effective to steer. Existing evaluations only find that steering at the LM head or embedding layer is ineffective (Subramani et al., 2022) and that steering both the attention and MLP blocks is more effective than only intervening on one of them (Zhang et al., 2024b).
@@
Should only one or many layers be controlled? - Many layers. It is generally assumed that intervening on all layers at once leads to more effective steering at a higher cost in the general capabilities of the model. However, to the best of our knowledge, there are no experiments proving this. Notably, many concepts can be most effectively steered in middle layers (Panickssery et al., 2024; Im & Li, 2025).
@@ 8.2 Commonly Controlled Concepts
{mdframed}
[style=takeawaybox] Takeaway: What concepts have been controlled with RepE? There are different types of concepts that can be controlled with RepE, among which high-level behavioral concepts like Truthfulness, Harmfulness, and Fairness are the most popular.
@@ 9.8.2 Games
{mdframed}
[style=takeawaybox] Takeaway: Applications RepE has been used in
  • AI Safety to reduce harmfulness, increase truthfulness, and control refusal behavior.
  • Interpretability to confirm the presence and shape of concept representations.
  • AI Ethics to reduce societal bias, align LLM behavior with human preferences, and control ethical reasoning.
  • Knowledge Editing to change or unlearn knowledge and adapt model encoded beliefs.
  • Other areas such as controlling task execution and text generation and improving general model performance.
@@ 10 Comparing RepE to Other Methods
10.1 Feature Identification and Steering
There is a range of methods that aim to identify activations in the model related to a target concept, which can then be used to steer the LLM’s internal processes or output behavior.
@@ 10.1 Feature Identification and Steering
Model Editing (ME). ME aims to make targeted modifications to the model weights that alter specific knowledge or behavior while preserving overall performance (Meng et al., 2022; 2023; Fang et al., 2024). After identifying relevant parameters, one can make precise changes and verify their impact on the model’s outputs. In contrast to RepE, ME does not directly attempt to identify representations in the activations space and instead only finds tweaks to the model parameters that bring the desired change.
Probing. In Probing, a classifier is trained to predict the occurrence of a concept in the model’s activations. The derived classifier can also be used to adapt the activations during inference. Section 6.1.2 explains that a probe can be used to dynamically decide the strength of the steering function. Furthermore, high scores of a probe on safety-relevant concepts, like task drift (Abdelnabi et al., 2024a) or deceptive reasoning (MacDiarmid et al., 2024; Goldowsky-Dill et al., 2025), can trigger a shutdown of the model or engage a more expensive safety procedure. While probing can be part of Representation Identification, it is usually not designed to exert control on model behavior.
Causal Abstraction (CA). In CA, a causal model for a task is specified, and an attempt is made to align the internal activations of a model with this high-level model (Geiger et al., 2024a; b). If intervening on the causal model has similar effects as intervening on the network representations, it can be said to be a good abstraction. In principle, this can allow us to identify and manipulate model representations. The focus of this method lies on its applications to interpretability, while RepE is also applied to control the LLM’s behavior.
Concept Erasure (CE). CE attempts to identify and remove a concept feature from a model’s representations. This requires identifying where the feature is represented and then removing it from the activations or weights. Indeed, some RepE methods attempt to erase concepts from the model representations (Li et al., 2024b; Rozanova et al., 2023).
@@ 10.2 Meta-study comparing to Other Methods For Behavior Control
Other methods for controlling the behavior of LLMs include prompting, fine-tuning, and decoding-based methods. We describe them, contrast them to RepE, and conduct a meta-survey that compares them to RepE according to their effectiveness and impacts on the model’s capabilities.

### v4

@@ 7.2.2 What Works for Representation Operationalization?
Are matrices or vectors better concept operators? - Matrix more effective but more costly.
Multiple works that propose using matrices instead of vectors as concept operators find that they improve steering effectiveness (Postmus & Abreu, 2024; Rajendran et al., 2024; Pham & Nguyen, 2024). Results from Zou et al. (2023a) are more ambivalent, finding that a matrix slightly outperforms a vector-based concept operator, but is clearly outperformed by a vector that is derived from additional forward passes for every new input. While matrices can be more effective, they can also come at an increased cost (Postmus & Abreu, 2024) or reduced output quality (Rajendran et al., 2024).
Are linear or non-linear concept operators better? - Non-linear.
While most methods perform linear operations on the representations, the direct comparison between linear and non-linear methods favors the non-linear ones. Non-linear Spectral Editing of Activations (SEA) outperforms linear SEA (Qiu et al., 2024), and non-linear ITI is more effective than ITI, although this comes at a higher cost in capabilities (Hoscilowicz et al., 2024).
@@ 7.2.3 What works for Representation Control?
Is it more effective to modify weights or activations? - Unclear.
@@
This question is still open. Wang et al. (2024a) find that their method for controlling model weights is more effective than intervening on activations while leading to a larger reduction in output quality. However, a broader evaluation of that question is necessary.
Which steering function should be used for modifying activations? - Unclear.
Zou et al. (2023a)
@@
find that a piece-wise operation is more effective than linear addition. Chu et al. (2024) find that projection outperforms a product, which again is more effective than linear addition. Luo et al. (2024) find that linear addition is more effective than orthogonal projections at a higher cost in capabilities. Linear addition remains the most popular activation steering function despite some other steering functions outperforming it. This warrants further investigation comparing different operations. Krasheninnikov & Krueger (2024) study different activation steering functions in a toy classification setup. They find that more expressive steering functions have a higher performance ceiling and can control more complex tasks, but simpler steering functions are effective with lower amounts of data.
At which token positions can activations be steered most effectively? - Apply steering at every inference step.
@@
Overall, it is more effective to continuously apply RepE at every inference step than to only apply it at the first token (Subramani et al., 2022). This is because the steering effect diminishes throughout the generation (Scalena et al., 2024). However, only intervening on the first token can still be very effective (Subramani et al., 2022) and has a lower impact on output quality (Scalena et al., 2024).
Which component of the transformer should be steered? - Unclear.
@@
RepE often modifies activations in the residual stream, the attention or the MLP block, but it is unclear which component is most effective to steer. Existing evaluations only find that steering at the LM head or embedding layer is ineffective (Subramani et al., 2022) and that steering both the attention and MLP blocks is more effective than only intervening on one of them (Zhang et al., 2024b).
Should only one or many layers be controlled? - Many layers.
@@
It is generally assumed that intervening on all layers at once leads to more effective steering at a higher cost in the general capabilities of the model. However, to the best of our knowledge, there are no experiments proving this. Notably, many concepts can be most effectively steered in middle layers (Panickssery et al., 2024; Im & Li, 2025).
8 Evaluation of RepE Methods
Good practices and resources for evaluations are a key drivers of progress for any area of Machine Learning. However, for RepE evaluations are not standardized and metrics are still developing. This section outlines current and best practices for evaluating RepE methods and describes existing and potential future benchmarks.
8.1 Common Evaluation Methodologies
The general setup for evaluating RepE methods involves measuring the steering effect, a measure of how much the behavior of model with regard to the concept of interest was changed. For this we use the model before and after intervention to generate outputs on a dataset or task that is often related to the concept. For the generated outputs we use a metric that determines in how much the outputs are aligned with the concept. Additionally it’s common to evaluate in how much the general model quality deteriorates.
8.1.1 Evaluation Task and Dataset
Depending on the concept of interest, different tasks and datasets are used. It is common to use established datasets and benchmarks that cover the steered concept or targeted downstream application. Commonly used datasets for evaluating changes in a concept are TruthfulQA (Lin et al., 2022) for increasing truthfulness, AdvBench (Zou et al., 2023b) for preventing harmful outputs or BOLD (Dhamala et al., 2021) for reducing societal bias (see Section 10 for more common datasets per concept). Other papers devise their own specific tests and dataset, for example to measure causal efficacy of an identified concept operator (Arora et al., 2024), to evaluate steering effect a wide range of concepts (Wu et al., 2025) or to study RepE in a toy setting (Krasheninnikov & Krueger, 2024).
A key difference between evaluations is whether the model is asked to generate answers to multiple-choice questions or make longer generations. While evaluations on multiple-choice questions are convenient to run, they are less faithful to downstream settings in which a model will be employed (Pres et al., 2024).
8.1.2 Measuring Steerability
To evaluate the steering effect, it is necessary to measure in how much a model’s outputs are aligned with the concept. This requires a way to label outputs depending on their alignment with the concept and to construct a metric out of this. A key difference in evaluations of the steering effect is if the model gets to freely generate any output that then gets evaluated or whether we are evaluating the likelihood of the model generating specific pre-defined answers.
Table 9: Source of labels and metrics for evaluating LLMs on different generation
Form of generation Free generation Likelihood on specified answers
Source of Labels LLM-judge, text classifier, word occurrences correct answers, ground-truth text
Metrics Judge Score, # of concept-related words Accuracy, Logprob-based metrics
Source of Labels Evaluation methodologies differ on how they determine the amount of concept alignment of an output. Sources of labels can be LLM judges, ground-truth labels or text of a datasets, and word occurrences in generated text.
LLM judges are commonly used to evaluate freely generated text. Many papers instruct a frontier LLM such as GPT-4 to give a numerical judgment about how much a generated text aligns with the concept of interest (Wu et al., 2025). Other papers use smaller models that are specialized for judging a specific property and have been validated for this task. For example Wang & Shu (2023) use small, specialised LLM for toxicity classification (Caselli et al., 2021) and a lexicon and rule-based tool for sentiment scoring (Hutto, 2020).
Many papers use exiting labels in a dataset. In multiple-choice benchmarks such as MMLU, the correct labels can be used as ground truth to judge the models performance. Furthermore, on a texts related to the concept, the ability to correctly predict the correct next token can be taken as a source of information about the concept alignment of a model (van der Weij et al., 2024). Other papers judge outputs based on how often specific concept-related words occur. Human judgments are rarely used, since they are expensive. Only Wang et al. (2024c) use a small set of human evaluators to confirm that their automated metrics align with human judgment.
Metrics for the Steering Effect Depending whether pre-specified or freely generated outputs are used, different metrics for the steering effect are available.
Evaluation methodologies that use a judge for scoring usually simply take the score assigned by the judge as their metric. For example LLM judges are often instructed to provide a score (e.g. from 0-4), while specialized judging-models might directly output a concept score.
Some concepts can also be measured by detecting or counting concept-related words and phrases in freely generated outputs. A common technique for measuring the Attack Success Rate of jailbreaks is to measure the look the presence of specific words and phrases such as "I am sorry as an LLM…" (Arditi et al., 2024). Turner et al. (2024) count the number of wedding-related words and Jorgensen et al. (2024) count the number of words related to specific words to measure whether they were able to shift the topic and genre of generated text.
On multiple-choice questions many papers measure the average accuracy. On longer generations, van der Weij et al. (2024) measure the difference in the top-1 accuracy at predicting the correct next token on datasets that are or aren’t related to the concept. However, this ignores shifts in the probability the model assigns to the correct answer. Thus, Panickssery et al. (2024) evaluate steering effect as the average likelihood the model assigns to the correct response. Tan et al. (2024) instead use the logit-difference propensity, which is calculated as the difference in logits between the positive and negative answer. They argue that excluding the normalization provided by the last softmax makes the metric more linear with respect to the model’s activations. Alternatively, Turner et al. (2024) measure the probability of the correct answer being in the top-k tokens. Arora et al. (2024) devise a log-odds metric that measures the causal efficacy of an intervention on the model’s representation. Krasheninnikov & Krueger (2024) count the fraction of examples where the model can change the desired attribute without changing any other attributes. Finally, Pres et al. (2024) take the difference in log-likelihoods for positive and negative continuations, normalize them and then take the mean difference between the base and steered model.
8.1.3 Measuring Changes in Model Quality
Interventions on the models representations should have minimal impact on the general quality and capabilities of the model. RepE papers use a variety of measures to evaluate their impact on model quality.
For this, many papers report scores on popular capabilities benchmarks like MMLU (Hendrycks et al., 2021) or MT-Bench (Zheng et al., 2023a) as a measure of models quality. Other papers evaluate the models capability to correctly predict the next token by running it on a general dataset like the Pile (Gao et al., 2020) and evaluating Perplexity(Scalena et al., 2024), top-1 accuracy (van der Weij et al., 2024) or the probability to have the correct token in the top-k tokens (Turner et al., 2024). Furthermore, it is possible to evaluate specific aspects of model performance that could be deteriorated by RepE. Brumley et al. (2024) measure diversity of generations bi- and tri-gram entropies. Wu et al. (2025) evaluate the fluency and instruction following of the model via an LLM-judge.
Lastly, some papers decide to combine steering effectiveness and model quality into one performance metric for RepE. von Rütte et al. (2024) propose perplexity-normalized effect size, which divides the increase in probability of the concept being present by the decrease in model fluency. Wu et al. (2025) report the harmonic mean of the score for concept-alignment, instruction following and fluency as a combined metric.
8.1.4 Other Experimental Details
Hyperparameter Ablations: RepE methods are commonly evaluated on a range of hyperparameters and settings. For many RepE methods the steering strength is a hyperparameter that guides how strong the intervention is. Altering the steering strength can outline the tradeoff between steering effect and model quality. Furthermore, this indicates the quality of the concept operator is, since increasing the steering factor should lead to a monotonic or even linear increase in concept-alignment (Pres et al., 2024; Tan et al., 2024). Other common hyperparameters to be altered are the sets of layers on which the intervention is conducted, the token position to intervene on, the number of samples used in Representation Identification and the family and size of models.
Baselines: To measure the steering effect and reduction in model quality the steered model must be compared to the original model without intervention. However, we also need to compare to other control models. It’s common to compare to earlier RepE methods. Furthermore, many papers make comparisons to prompting or fine-tuning methods.
8.2 Best Practices for Evaluating RepE Methods
Free generation vs likelihood of specified answers. Measuring the steering through measures that use pre-specified answers has the advantage that one can measure shifts in the log-likelihood of the model. Pres et al. (2024) argue this is important because disregarding the confidences of the model, loses information about how variable behavioral expressions are. On the other hand, when none of the pre-specified answers has a high likelihood, such metrics would not take into account the actual generations a model would produce thus miss important aspects of its behavior.
Using specialized judges. Many studies rely on LLM-as-a-judge setups, where a highly capable and general model is instructed to rate generated text. However, using LLMs-as-a-judge is known to be biased, sometimes inaccurate and is usually not scientifically validated as a metric (Fu et al., 2024; Zheng et al., 2023b). When possible it might be preferrable to use specialised models that have been developed and validated to score a specific concept, such as small LLMs trained for detecting specific concepts (Inan et al., 2023; Caselli et al., 2021; Dementieva et al., 2023).
Varying Steering Strength. In many RepE methods it is possible to modulate the strength of the intervention. By showing the steering effect and model quality at different steering strengths readers can get a sense of the tradeoff between these properties. Furthermore, it is useful to plot the relationship between steering strength and steering effect. If the steering effect monotonically, linearly or strongly increases with higher steering strength this showcases that the RepE can effectively modulate the desired concept. Tan et al. (2024) even propose a metric based on this relationship. For the same reason it is informative to show the effect if a negative steering strength is applied.
Realistic and Difficult Test Settings. While it is convenient to use multiple-choice format to evaluate the steering effect, Pres et al. (2024) argue that it is important to the effect on open-ended generations since this setting is more similar to downstream settings where RepE will be applied. Additionally, it is preferable to evaluate RepE on difficult settings that are similar to real world deployment. For example, when using RepE to defend against jailbreaks it is important to use state-of-the-art attacks and evaluate multi-turn jailbreaks.
Reporting Changes in Models’ Quality. Only a minority ( 35%) of papers report the change in models’ capabilities after applying Representation Engineering. Since we know that RepE can reduce these capabilities, it is important to measure the LLMs’ quality to compare the merits of each method. Furthermore, it is desirable to administer diverse capability tests, like measuring the ability to generate fluent text while also the evaluating the accuracy at answering multiple-choice questions.
Measuring Generalization. RepE methods can struggle to generalize to different settings (see Section 13.1). Thus, it is not sufficient to evaluate the steering effect on another subset of the dataset used for identifying the concept operator. Instead studies should systematically vary aspects of the settings between Representation Identification and Control.
Comparing to Strong Baselines. When proposing a new method, its effectiveness should be compared to the current state-of-the-art RepE method. However, often papers only compare to early methods that have since been surpassed. Similarly, when comparing to other families of methods, studies should not compare to a naive prompting or fine-tuning approach, but to the state-of-the-art method for their respective problem.
Evaluating Diverse Concepts. The effectiveness of RepE methods should be measured on a diverse range of concepts. Since steerability can differ between concepts it is difficult to assess the quality of a method on only one concept. Furthermore, it is desirable to evaluate on different types of concepts (see Section 9.1) to determine whether a method has strengths or weaknesses at steering some type of concept.
Testing on Different Model Sizes. Appendix A.1 finds that most RepE experiments are conducted on models between 3-10 billion parameters. To prove the general effectiveness of a RepE method, it is necessary to evaluate it on models of different sizes. Specifically, larger models need to be tested to indicate that the method scales to frontier models.
8.3 RepE benchmarks
Existing Benchmarks. There are two benchmarks that attempt to evaluate and compare RepE methods (Wu et al., 2025; Im & Li, 2025).
AxBench (Wu et al., 2025) is a benchmark that evaluates the ability of RepE methods to identify concept operators that accurately detect the presence of a concept and steer the LLMs’ behaviour. For this, they built a synthetic dataset of positive and negative examples for 500 concepts. Using this dataset, 9 RepE methods are used to derive concept operators and compared to prompting and fine-tuning. Steering ability is determined by generating responses to instructions while performing Representation Control and then using an LLM to score the generation on presence of the concept, instruction following and fluency of the generation. For concept detection Difference-in-Means, Logistic Regression and ReFT-r1 perform best. However, prompting outperforms all RepE methods at steering LLM behaviour to a concept. Notably, all tested concepts are of a similar type, testing whether a concrete, narrow topic is contained in a text. This benchmark could be extended by consider different types of concepts.
Im & Li (2025) evaluate four methods for Representation Identification on their ability to steer outputs positively and negatively towards a concept on multiple-choice and open-ended generation tasks. They find that Difference-in-Means outperforms other methods to identify a steering vector. This benchmark could be extended by comparing different methods for Representation Operationalization or Control and evaluating them on a more diverse range of concepts.
Desireable Benchmark Properties. There is currently no benchmark that provides a way to comprehensively evaluate the effectiveness of a RepE pipeline. Thus we are currently unsure which pipeline is most effective in which situations. A benchmark can also serve as a guide for the future development of RepE methods by providing clarity about the effectiveness of methods and being a target to aim for. Such unified evaluations can also surface previously undiscovered limitations of methods (Brumley et al., 2024).
An ideal benchmark for RepE should cover a range of concepts of different types since different RepE methods might steer different concepts more effectively (Brumley et al., 2024). Hereby, the focus should lie on use cases where RepE is foreseen to be practically useful, such as improving truthfulness, changing goals, adapting situational awareness, or removing societal biases. Furthermore, evaluations should cover multiple models, test the influence of dataset size, and show the impact of different steering strengths. Additionally, it should specifically evaluate the weaknesses of current RepE methods (see Section 13.1), thus putting a spotlight on improvements on these challenges. Of course, this benchmark should make sure to follow best practices outlined above (Section 8.2).
@@ 9.2 Commonly Controlled Concepts
Takeaway: What concepts have been controlled with RepE?
There are different types of concepts that can be controlled with RepE, among which high-level behavioral concepts like Truthfulness, Harmfulness, and Fairness are the most popular.
@@ 10.8.2 Games
Takeaway: Applications
RepE has been used in
AI Safety to reduce harmfulness, increase truthfulness, and control refusal behavior. Interpretability to confirm the presence and shape of concept representations. AI Ethics to reduce societal bias, align LLM behavior with human preferences, and control ethical reasoning. Knowledge Editing to change or unlearn knowledge and adapt model encoded beliefs. Other areas such as controlling task execution and text generation and improving general model performance.
@@ 11 Comparing RepE to Other Methods
Table 12: Difference and Similarities of related Families of Methods to RepE
Other Method Difference to RepE
Prompting Changes inputs instead of activations or weights
Soft-prompting Operates on token embeddings instead of hidden states
Fine-tuning Doesn’t target specific concept representations
Decoding-based Methods Operates on logits instead of hidden states
Mechanistic Interpretability Bottom-up vs Top-Down view
Activation Patching Fully replaces activations instead of adapting them
Probing Only does Representation Identification and not Control
11.1 Related Methods
There is a range of related methods that attempt to identify how concepts are represented in Neural Networks or that aim to control the behavior of LLMs.
Prompting. Changing the (system-)prompt given to an LLM is the most straightforward and commonly used way to steer the behavior of a model. However, it does so without any changes to weights or activations and does not provide any interpretability benefits. Furthermore, RepE does not take up space in the context window, while the additional tokens from prompting, especially when using many in-context examples, increases the computational cost.
Soft-Prompting. Similar to RepE, soft-prompting modifies the model in a continuous embedding space. It does so by operating on the input embeddings, whereas RepE modifies internal activations or weights. However, soft-prompting does not have the goal of identifying the representation of a concept.
Fine-tuning. Fine-tuning modifies the weights of a pre-trained model by further training it on a specific dataset or reward signal. While fine-tuning does influence the model’s representations, it does not do so in a manner that is targeted to identifying and controlling representations of specific concepts. Instead, it applies some uninterpretable, non-sparse changes to the weights that lead to the desired outputs. Unlike RepE, fine-tuning does not allow for precise and isolated control of concepts of interest. Furthermore, many RepE methods provide continuous control, where a concept can be increased or decreased to the desired intensity. As mentioned in Section 6.2.1, there is an overlap between fine-tuning and RepE methods that do output optimization and apply a weight steering function.
Decoding-based methods. Decoding-based approaches steer the behavior of LLMs by shifting the token probabilities during the decoding step. Contrastive Decoding takes two sets of token probabilities, contrasts them against each other, and consequently shifts the token probabilities (Li et al., 2023b). Often, this means taking predictions from two different models, one of which is better at the desired behavior. Alternatively, contrastive predictions can be taken from the same model on contrastive inputs. Similar to RepE methods, this method steers the behavior of the model during inference time. However, contrastive decoding requires multiple forward passes per generated token. Furthermore, it only superficially shifts token probabilities and does not tap into the internal representations of LLMs.
@@ 11.1 Related Methods
Probing. In Probing, a classifier is trained to predict the occurrence of a concept in the model’s activations. Classically, probes are trained to interpret Neural Network. Furthermore, they can be used to detect concepts like task drift (Abdelnabi et al., 2024a) or deceptive reasoning (MacDiarmid et al., 2024; Goldowsky-Dill et al., 2025) during inference such that a more expensive safety procedure can be triggered. However, they can also be used as a method for Representation Identification in RepE or as a way to dynamically modulate the strength of the steering function.
@@ 11.2 Meta-study comparing to Other Methods For Behavior Control
Other methods for controlling the behavior of LLMs include prompting, fine-tuning, and decoding-based methods. We describe conduct a meta-survey that compares them to RepE according to their effectiveness and impacts on the model’s capabilities.
@@
A list of all papers from this metastudy is provided in D and Table 14 and Table 15 respectively show experimental results for each comparison of RepE with prompting and fine-tuning respectively.